OpenStudy (anonymous):

I give medals! Suppose a culture of bacteria starts with 8000 bacteria. After one hour, the count is 9000. Find an exponential equation that models the number of bacteria in hours. Keep at least 4 decimal places in your formula for rounded values. Find the doubling time of the bacteria. Round the doubling time to the nearest 0.01 hours.

OpenStudy (anonymous):

simple way or hard way?

OpenStudy (anonymous):

@satellite73 Whatever way you feel would be easiest to explain.

OpenStudy (anonymous):

in one hour it goes from 8000 to 9000 so it increases by a factor of \[\frac{9000}{8000}=\frac{9}{8}\] and therefore you can model it as \[8000\times \left(\frac{9}{8}\right)^{t}\] where \(t\) is time in hours pretty simple right?

OpenStudy (anonymous):

if you need to make your teacher happy and use \(8000e^{kt}\) we can do that too, but it takes much longer

OpenStudy (anonymous):

\[P(t)=P _{0}e ^{k*t}\] this is the correct equation right?

OpenStudy (anonymous):

ok damn i guess we have to use that i was hoping to get away with \(A=8000\times\left(\frac{9}{8}\right)^t\) but i guess not

OpenStudy (anonymous):

it isn't that bad, just takes a bit longer to do ready?

OpenStudy (anonymous):

@satellite73 Yeah, that's what I was thinking earlier but I confused myself :o

OpenStudy (anonymous):

ok so we know it is \[8000e^{kt}\] we just don't know \(k\)

OpenStudy (anonymous):

here is how we find it \[A=8000e^{kt}\] if \(t=1\) then \(A=9000\) so we have to solve \[9000=8000e^k\]for \(k\)

OpenStudy (anonymous):

first divide by 8000 and get \[\frac{9}{8}=e^k\] then take the log to get \[k=\ln(\frac{9}{8})\] then use a calculator to find this number

OpenStudy (anonymous):

I got 0.1178

OpenStudy (anonymous):

So, the equation would be \[P(t)= 8,000e ^{0.1178t}\] right?! :)

OpenStudy (anonymous):

let me check, looks good

OpenStudy (anonymous):

yeah that is right

OpenStudy (anonymous):

Alrighty! :) And for the second half I would just put into 2 for t right? Since that is doubled the time?

OpenStudy (anonymous):

oh no \(t\) is time, so if you put \(t=2\) you find out how many you have in two hours

OpenStudy (anonymous):

Ah! I misread the problem! :o

OpenStudy (anonymous):

to find the doubling time set \[\large e^{.1187t}=2\] and solve for \(t\)

OpenStudy (anonymous):

It's been a really long day! But thanks! I'll solve tat right now!

OpenStudy (anonymous):

kk i'll check if you like

OpenStudy (anonymous):

Please! I'll let you know when I am done

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[2=e ^{.1187*t}\] \[\frac{ \ln2 }{ 0.1187 }=\frac{ 0.1187t }{ 0.1187 }\] \[t=5.84 hours\]

OpenStudy (anonymous):

that is what i got too http://www.wolframalpha.com/input/?i=ln%282%29%2F.1187

OpenStudy (anonymous):

Thanks so much! I really appreciate it! My brain is too fried from this week!

OpenStudy (anonymous):

finals soon?

OpenStudy (anonymous):

oh, and your welcome, good luck with this, looks like you are doing fine

OpenStudy (anonymous):

Yeah, lots of them! Thanks good luck if you have any too! :)